OpenStudy (unklerhaukus):

Solve the autonomous differential equation. (independent variable is absent) \[\frac{\mathrm d^2y}{\mathrm dx^2}+n^2y=0\]

OpenStudy (sirm3d):

\[y=A\cos nx + B\sin nx\]

OpenStudy (unklerhaukus):

im having trouble with the autonomus method

OpenStudy (anonymous):

Is \(n\) just a constant?

OpenStudy (unklerhaukus):

yes

OpenStudy (anonymous):

This is the way I was taught to do these. Let \(u=\frac{dy}{dx}\). Then \[ \frac{d^2y}{dx^2} = \frac{du}{dx} = \frac{du}{dy} \frac{dy}{dx} = y\frac{du}{dy} \]and when you substitute than into the equation you get a first order ODE for \(u\) in terms of \(y\). Solve for \(u\).

OpenStudy (anonymous):

Sorry, that should be \(u\frac{du}{dy}\) at the end there. I knew something didn't look right...

OpenStudy (anonymous):

Possibly a better way to do it is to assume that \(e^{\lambda x}\) is a solution, which if you substitute into the equation, gives \[\lambda^2 e^{\lambda x} + n^2e^{\lambda x} = 0 \]and of course you can divide out \(e^{\lambda x}\), so \[\lambda^2 + n^2 = 0 \]so \(\lambda = \pm in\), and so your general solution is \[ y=ae^{in} + be^{-in} \]where \(a\) and \(b\) are arbitrary constants. You can put that in trig form using Euler's identity.

OpenStudy (unklerhaukus):

\[\newcommand \p \newcommand \p \dd [1] {\,\mathrm d#1 } \p \de [2] {\frac{\mathrm d #1}{\mathrm d#2} } \p \den [3] {\frac{\mathrm d^#3 #1}{\mathrm d#2^#3}} \begin{align} \den yx2+n^2y&=0 &\text{$x$-absent} \\ &&\text {let }\de yx&=u \\ &&\den yx2&=\de ux=\de uy\de yx=\de uyu \\ \de uyu+n^2y &=0 \\ u\de uy &=-n^2y \\ u\dd u &=-n^2y\dd y \\ \int u\dd uc &=-n^2\int y\dd y \\ \frac{u^2}2 &=-\frac{n^2y^2+c}2 \\ u^2 &=-n^2y^2+c \\ % \de yx &=\sqrt{} % &=a\cos nx+b\sin nx \end{align}\]

OpenStudy (anonymous):

\[\frac{ d ^{2}y }{ dx ^{2} }+n ^{2}y=0\rightarrow \frac{ d ^{2}y }{ -n ^{2}y }=dx ^{2}\rightarrow \int\limits_{}^{}\int\limits_{}^{}\frac{ d ^{2y} }{ -n ^{2}y }=\int\limits_{}^{}\int\limits_{}^{}dx ^{2}\rightarrow -\frac{ 1 }{ n ^{2} }(y \ln (y)-y)=\frac{ x ^{2} }{ 2 }+c _{1}x-c _{2}\] figure out the y yourself

OpenStudy (unklerhaukus):

\[ \newcommand \p \newcommand \p \dd [1] {\,\mathrm d#1 } \p \de [2] {\frac{\mathrm d #1}{\mathrm d#2} } \p \den [3] {\frac{\mathrm d^#3 #1}{\mathrm d#2^#3} } \p \pa [2] {\frac{\partial #1}{\partial #2} } \p \pan [3] {\frac{\partial^#3 #1}{\partial#2^#3} } \p \pat [2] {\tfrac{\partial #1}{\partial #2} } \begin{align} \den yx2+n^2y&=0 &\text{$x$-absent} \\ &&\text {let }\de yx &=p \\ &&\den yx2 &=\de px=\de py\de yx=p\de py \\ p\de py+n^2y &=0 \\ p\de py &=-n^2y \\ p\dd p &=-n^2y\dd y \\ \int p\dd p &=-n^2\int y\dd y \\ \frac{p^2}2&=\frac{-n^2y^2+c}2 \\ p^2 &=n^2(a^2-y^2) \\ p &=n\sqrt{a^2-y^2} \\ \de yx &=n\sqrt{a^2-y^2} \\ \frac{\dd y}{\sqrt{a^2-y^2}} &=n\dd x \\ \int\frac{\dd y}{\sqrt{a^2-y^2}} &=n\int\dd x \\ \arcsin \Big(\frac ya\Big) &=nx+b \\ y &=a\sin(nx+b) \\ y &=a\big(\sin(nx)\cos b +\cos(nx)\sin b\big) \\ \\ y(x) &=A\sin (nx)+B\cos (nx) \end{align}\]

OpenStudy (unklerhaukus):

That bit that goes off the edge ***\[\newcommand \p \newcommand \p \dd [1] {\,\mathrm d#1 } \p \de [2] {\frac{\mathrm d #1}{\mathrm d#2} } \begin{align} && &\text{$x$-absent} \\ &&\text {let }\de yx &=p \\ &&\den yx2 &=\de px=\de py\de yx=p\de py \\ \end{align}\]

OpenStudy (anonymous):

u could take it as a separable equation(the easiest way!)

OpenStudy (unklerhaukus):

seperable, but you have to take double integral ,

OpenStudy (anonymous):

but still the easiest way

OpenStudy (unklerhaukus):

dx^2 is diffferent to d^2x and i get confused, i havent learnt that technique in class yet

OpenStudy (anonymous):

it's just how u write it ! u must understand how they're different

OpenStudy (unklerhaukus):

it think the eaisiest way is to take the charestic equation m^2+n^2=0 m=±in y=Acos nx+Bsin nx but that wasnt the technique i was trying

OpenStudy (anonymous):

and also u could take it as a constant coefficient equation

OpenStudy (unklerhaukus):

this is a copy of your method (that dosent run off the page)\[\frac{ d ^{2}y }{ dx ^{2} }+n ^{2}y=0\\\rightarrow \frac{ d ^{2}y }{ -n ^{2}y }=dx ^{2}\\\rightarrow \int\limits_{}^{}\int\limits_{}^{}\frac{ d ^{2y} }{ -n ^{2}y }=\int\limits_{}^{}\int\limits_{}^{}dx ^{2}\\\rightarrow -\frac{ 1 }{ n ^{2} }(y \ln (y)-y)\\=\frac{ x ^{2} }{ 2 }+c _{1}x-c _{2}\]

OpenStudy (anonymous):

your technique is for complicated combination of y' and y"

OpenStudy (unklerhaukus):

x- absent and y- absent are thte first two types of DE's in the chapter on 2nd order DEs this question was from the first problem set

OpenStudy (anonymous):

lots of ways u can solve a differential equation but u must choose the best one. u see I'm getting ready for my college entrance which is the hardest problems of differential equations will be presented in the exam that is why I was insisting on separable!

OpenStudy (unklerhaukus):

Yes, your right seperation of variables has got to be the most elementry technique,

OpenStudy (unklerhaukus):

i just haven't learnt it properlly in my class, what is the difference between \[\iint \mathrm d^2x\] and \[\iint \mathrm dz^2\]

OpenStudy (anonymous):

what do u think of this ? what kind of equation is this?

OpenStudy (unklerhaukus):

\[\sin y\frac{\mathrm dy}{\mathrm dx}=\cos y (1-x\cos y)\]

OpenStudy (anonymous):

in dz2 z must be the dependent variable otherwise u can't take double integral! u know what I mean?

OpenStudy (unklerhaukus):

okay like i should of used y

OpenStudy (unklerhaukus):

go on

OpenStudy (anonymous):

\[d(x ^{2}) is different from dx ^{2}\]

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

\[d(x ^{2})\] is different from \[dx ^{2}\]

OpenStudy (anonymous):

u get it?

OpenStudy (unklerhaukus):

ok so i should have asked the difference between \(\iint \mathrm d^2x\) and \(\iint \mathrm d(y^2)\)

OpenStudy (anonymous):

\[d ^{2}x =(dx)(dx)\] but in \[∬d(y ^{2}) \] you're integrating with respect to \[y\]

OpenStudy (anonymous):

sorry that is \[y ^{2}\]

OpenStudy (anonymous):

you're a college student ? right? what major?

OpenStudy (unklerhaukus):

\[\sin y\frac{\mathrm dy}{\mathrm dx}=\cos y (1-x\cos y)\\ \cos y (x\cos y-1)\mathrm dx+\sin y\mathrm dy=0\\ \\\,\\ M_x=\cos^2y\qquad N_y=\cos y\qquad \text{inexact}\\ \qquad \,\\ \frac{N_y-M_x}{M}=\frac{\cos y(1-\cos y)}{\cos^2y}=\sec y-1\] Integrating factor \[R(y)=\int(\sec y-1)\mathrm d y\]

OpenStudy (unklerhaukus):

hmm its not so easy to that integral of sec y

OpenStudy (anonymous):

if u take v=sec(y) it would become a simple first order equation

OpenStudy (anonymous):

now just see how simple is that if u look closely to the equation?